# Homeostazis and Sparsity in Transformer

### v1

@@ Abstract
The transformer architecture has become an integral part of the field of modern neural networks, playing a crucial role in a variety of tasks, such as text generation, machine translation, image and audio processing, among others. There is also an alternative approach to building intelligent systems, proposed by Jeff Hawkins and inspired by the processes occurring in the neocortex. In our article we want to combine some of these ideas and to propose the use of homeostazis mechanisms, such as RFB-kWTA and "Smart" Inhibition, in the attention mechanism of the transformer and at the output of the transformer block, as well as conducting an experiment involving the introduction of sparse distributed representations of the transformer at various points. RFB-kWTA utilizes statistics of layer activations across time to adjust the entire layer, enhancing the values of rare activations while reducing those of frequent ones. "Smart" Inhibition also uses activation statistics to sample sparsity masks, with rarer activation times are more likely to be activated. Our proposed mechanisms significantly outperform the classical transformer 0.2768 BLEU and a model that only makes use of dropout in the attention mechanism and output of the transformer block 0.3007 BLEU, achieving a score of 0.3062 on the Multi30K dataset.
Keywords transformer  $\cdot$ sparse distributed representations  $\cdot$ homeostazis  $\cdot$ natural language processing  $\cdot$ machine translation
@@ 4 Discussion
    The RFB mechanism in combination with Dropout has a positive effect on the generalizing ability of the model. In order to allow "subtle" patterns to periodically appear among the signs, we introduced a mechanism of homeostazis based on activation statistics. This mechanism makes it possible to enhance signs that are rare in time, which may contain those very "subtle" patterns. However, substituting dropout with “smart” inhibition allows for further improvement in the quality of the model.
The mechanisms proposed by us, due to their self-regulation, outperform the classical transformer in the task of machine translation. At the same time, all models incorporating a homeostazis mechanism, even those with $s=0.5$ have proven to be superior to the classic transformer model.

### v2

# Homeostasis and Sparsity in Transformer

@@ Abstract
The transformer architecture has become an integral part of the field of modern neural networks, playing a crucial role in a variety of tasks, such as text generation, machine translation, image and audio processing, among others. There is also an alternative approach to building intelligent systems, proposed by Jeff Hawkins and inspired by the processes occurring in the neocortex. In our article we want to combine some of these ideas and to propose the use of homeostasis mechanisms, such as RFB-kWTA and "Smart" Inhibition, in the attention mechanism of the transformer and at the output of the transformer block, as well as conducting an experiment involving the introduction of sparse distributed representations of the transformer at various points. RFB-kWTA utilizes statistics of layer activations across time to adjust the entire layer, enhancing the values of rare activations while reducing those of frequent ones. "Smart" Inhibition also uses activation statistics to sample sparsity masks, with rarer activation times are more likely to be activated. Our proposed mechanisms significantly outperform the classical transformer 0.2768 BLEU and a model that only makes use of dropout in the attention mechanism and output of the transformer block 0.3007 BLEU, achieving a score of 0.3062 on the Multi30K dataset.
Keywords transformer  $\cdot$ sparse distributed representations  $\cdot$ homeostasis  $\cdot$ natural language processing  $\cdot$ machine translation
@@ 4 Discussion
    The RFB mechanism in combination with Dropout has a positive effect on the generalizing ability of the model. In order to allow "subtle" patterns to periodically appear among the signs, we introduced a mechanism of homeostasis based on activation statistics. This mechanism makes it possible to enhance signs that are rare in time, which may contain those very "subtle" patterns. However, substituting dropout with “smart” inhibition allows for further improvement in the quality of the model.
The mechanisms proposed by us, due to their self-regulation, outperform the classical transformer in the task of machine translation. At the same time, all models incorporating a homeostasis mechanism, even those with $s=0.5$ have proven to be superior to the classic transformer model.